OpenStudy (anonymous):

. What is the approximate volume of a sphere with radius of 11 inches?

OpenStudy (anonymous):

\[V = \frac{ 4 }{ 3 } \pi r ^{3}\] is the formula for the volume of a sphere \[V = \frac{ 4 }{ 3 } \pi 11 ^{3}\ \]